40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching snip_of_beam_profiling_6_27.JPGP1020518.JPG
    Reply  Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching 130703-095332.jpg
       Reply  Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching P1020690.JPGP1020692.JPG
          Reply  Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching P1020705.JPGP1020703.JPG7_11_13_beam_profile.PNG7_11_13_beam_from_fixed_lens.PNG
             Reply  Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
                Reply  Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 130805-120628.jpg
                   Reply  Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 7_16_13_determining_polarization_axis.PNGTEK0000.PNG
                      Reply  Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                         Reply  Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                            Reply  Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam shielded_10s_TEK0005.PNGunshielded_10s_TEK0027.PNG
                               Reply  Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations 
                               Reply  Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table P1030056.JPGP1030057.JPGP1030059.JPGP1030056.JPG
                                  Reply  Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                     Reply  Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                        Reply  Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets w_o_capacitor_7_26_13.JPGcapacitor_7_26_13.JPGP1030066.JPGP1030070.JPG
                                           Reply  Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal 
                                              Reply  Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam 731_7graph.PNG731_11graph.PNGP1030318.JPG
                                                 Reply  Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam chmbrtmp.JPGfringe.JPG
                                                    Reply  Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam P1030403.JPG
                                                       Reply  Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency P1030409.JPGP1030417.JPGP1030420.JPG
                                                          Reply  Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. noise_budget.PNGP1030427.JPG
                                                             Reply  Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise P1030458.JPGP1030459.JPGP1030460.JPG
                                                       Reply  Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment P1030461.JPGP1030454.JPGP1030456.JPGP1030457.JPGAD590circuit2.PNG
                                                          Reply  Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise 
                   Reply  Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination new_setup.PNGbeam_recombination2.PNG
                   Reply  Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber 130805-121449.jpg
    Reply  Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO 
Message ID: 1240     Entry time: Tue Jul 16 10:27:47 2013     In reply to: 1237     Reply to this: 1241
Author: Erica 
Type: Notes 
Category: fiber optic 
Subject: mode matching, beam recombination 
July 14, 2013

We just went with the setup from yesterday.

We adjusted the input fiber so that one axis is vertical. If the lens of the collimator is as close to the fiber as possible, then the two screws on the side of the collimator are perpendicular to the little notch on the collimator, which is supposed to line up with one of the axes. Then we moved around the other knobs so we ended up getting the same maximum power output

At the output, we aligned everything up. After pol. beam splitter, there was a lot of power coming out of both s and p faces. This was confusing because we used a polarizing beam splitter to pick it off from the main beam.

We checked the beam going into the fiber was linearly polarized with a polarizing beam splitter - which it was, s-polarized.

We checked the output beam - took out the 45s beam splitter and then put pol beam splitter in there, rotated fiber until s- polarization was at a max (reflected beam)


We replaced the 45s beam splitter, still a lot of power not coming out in the direction we wanted - got transmitted rather than reflected. However, the power output from each beam after the 45s BS was about equal so didn't do anything about it.


After we checked the polarizations, we placed the polarizing BS to look at the beam recombination, and there was still light transmitted and reflected. This could be that the beam was really messy, and since the polarizing beam splitter is not 100%, there was still some p-polarization. This seems to be the reason why Tara had me place this second polarizing BS.

We placed a lens (f=70mm) placed in front of output to focus the beam, profile more like that from the laser

Using a OD2 filter, we used a photodiode (PDA100A) and an oscilloscope but could not see anything. Can't see fringes, just a DC line. Translating stage doesn't do much, we fiddled around with the knobs on mirrors and stage but not much change

Then we switched to looking at the Gaussian with DataRay, but - Gaussian from beam, not Gaussian from laser - clipping edge of mirrors so had to realign.

Here we can see fringes but the beams aren't very aligned. If they were exactly aligned, then you would see total destructive and total constructive interference. Even with moving knobs, it didn't really improve. Will work it more tomorrow.





Note: The power output from both beams should be about equal intensity.
ELOG V3.1.3-